heygirl:

20 × 70 = 2 tens × 7 tens = hundreds

dude:

This question isnt clear, maybe take a SS?

heygirl:

20 × 70 = 2 tens × 7 tens =

dude:

20*70 is not equal to 2*7 Not sure what the question is

heygirl:

it dor lil sis i dont no how to do that bc i am in 7 th but thanks eneyway

CraftyType1:

20 x 70 = 2 tens x 7 tens = 140 tens(1400)

jhonyy9:

my idea 20 = 2 tens 70 = 7 tens 20 × 70 = 2 tens × 7 tens = 2×7×ten×ten = 14×ten×ten =140×10=1400 hope helped